A Complete Bibliography of Publications in Marine Policy (1970–1979)

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
             WWW URL: http://www.math.utah.edu/~beebe/

15 January 2022 Version 1.01

Title word cross-reference

$10.00 [Ogl77, Pri77b]. $18.15 [Arc78, Cla78]. $22.00 [She78]. $22.80 [Mei78]. $25.00 [Wat78]. $27.50 [Por77]. $48.50 [Tet78]. $49.00 [Daw78]. $7.50 [Bir77a]. $8.50 [Ano77a]. $9.50 [Cus79].

1 [Bro78b]. 12.60 [Cla79]. 123pp. [Nas79]. 13.50 [Pul77b]. 15.00 [Met78]. 170-250 [Ano78j]. 170-252 [Ano79g]. 178-272 [Ano77h]. 178pp. [O’C79]. 19 [Kim79a]. 1939 [O’C79]. 1945 [Wat79d]. 1976 [Abe77, Kim77a, Law77b, Pul77a, Sta77, Yon77, Col77a, Gul77b]. 1977 [Bar78, Bir77b, Hju77, Kau77, Kim77b, Lof78d, Low78, Sco78, Wat77b]. 1978 [Abr79, Len78]. 1978/79 [Len79]. 1979 [Ano79f, Gil79, Kau79, Kim79a, Sha79a].

2 [Bro78c, Sha79a]. 2.50 [Cas77]. 20.00 [Daw78, Len77b]. 200-mile [Bir78a, Chu77a]. 25 [Len79]. 25.55 [Sha79b]. 254-332 [Ano79i]. 254-348 [Ano78l]. 27.50 [Len78]. 274-364 [Ano77j]. 28 [Wat79d]. 2nd [Rus78].

3.00 [Rus78]. 30.00 [Ion77]. 30th [BS79].

51.00 [Mor78].

6.50 [Wat77a].

77 [Len77b, Len77a]. 78-167 [Ano79j].

86-168 [Ano78m].

9.00 [Pul79a]. 9.75 [Pri77a]. 9.95 [O’C77]. 90-176 [Ano77k].

abalone [WMCS77]. Academic [Mor78]. Academy [Nas79]. access [AH79]. accession [Pan79]. act [Kel78]. action [Sal78]. activities [You79]. adaptation [Ric79c]. adolescent [Ano77o]. African [Dar77]. against [Sha79a]. Agency [Law77b]. agreements [CS79]. Agriculture [Hju77, Law77b]. Alternatives [Gul78]. America [Pal79]. Amsterdam [Daw78]. analysis [Fri78, Met79, Sto79, Cla79]. anchoveta [Gla79]. Anderson [Mei78]. Ann [Cla78, Cla79, Mei78, Cla78, Cla79, Mei78]. annotated [Chu77a]. Annual [Bar78]. Antarctica [Mit77]. Anticipating [Fro77]. appraisal [Rus78]. approach [Fro77, Wat79a]. approaches [Sha79a]. April [Ano77k, Ano78m, Ano79f, Ano79j, Hju77, Kim79a]. Aquaculture [Pal79, Her78, Pul77c, Sak79, Pul77a, Nas79, She78]. Arbor [Cla78, Cla79, Mei78]. archaeology [Jon78]. Archipelagos [And78]. area [Pri78, Hol78]. Armstrong [Cla79]. Arne [Chu78b]. Arvid [Ogl77]. Asia [Buz79]. aspects [AR77, Jac79]. assembly [Sco78, Yon77]. Assessment [Bir78a]. Assistance [Law77b]. Atlantic [Ber77, Cus77, Mac79, TL78, Wat78]. attitudes [Arc77]. August [Gul77b, Kim77a]. auspices [Nas79]. Australian [Law77b]. Auswirkungen [Wat79c]. Avenue [Cus79].

B [Cas77, Low77a]. balanced [Cus79]. Ballinger [Arc78, She78]. Barrage [Rus78]. Basic [Has77]. basin [You78a]. Bell [She78]. Bertram [Wat78]. biennial [Low78]. bilateral [CS79]. Birmingham [Sta77]. Biscayne [Kau79]. Board [Ano77b, Ano78b, Ano79e]. Book [Abr79, Ano77a, Arc78, Bir77a, Bir78a, Cas77, Chu77b, Chu78a, Chu78b, Cla78, Cla79, Cus79, Daw78, Ion77, Len77b, Len77a, Len78, Len79, Low77a, Mei78, Met78, Met79, Mor78, Nas79, O’C77, O’C78, O’C79, Ogl77, Ogl79, Por77, Pri77a, Pri77b, Pul77b, Pul79a, Rus78, Sha79b, She78, Tet78, Wat77a, Wat78, Wat79a, Wat79b, Wat79c, You77, You78b, You78a]. Books [Pul77b, Pul79a, Wat77a, Wat79a, You78b, You78a]. Booth [O’C77]. Bordeaux [Lof78d]. Borgese [Ogl77, Ogl79]. boundaries [Bro78a]. boundary [Gau79]. Bristol [Rus78]. British [O’C78, O’C79]. Brook [Met79]. Brooklyn [Cus79]. Bryan [O’C79]. bureaucratic [Fri78]. Burke [Bir77a].

C [Cla79, Met78]. Calendar [Ano79a, Ano79b, Ano79c, Ano79d]. California [WMCS77]. Cambridge [Arc78, Met78, She78]. Canada [Kau77, Mac79]. Canary [Ric79a]. Canterbery [She78]. Carl [Chu78a]. carriers [Man79]. case [Bro78a, Sha79a, Wat79b]. Center [Sno78]. Central [Abr79, Low77a]. Centralized [Kin79]. challenge [Kie77]. change [O’C79]. changes [Par79]. changing [You78a]. charting [Has77]. China [Val78]. Christoph [Wat78]. Churchill [O’C78]. cities [Kie77, Woo78]. Civil [Rus78]. claims [Chu77a]. Clark [Tet78]. Cliffs [Sha79b]. ClNECA [Tom79]. Closing [Lof78b]. coast [Mac79]. Coastal [Pul77c, Sno78, Car78a, Her78, Sak79, Tet78, You77, Cla78, Tet78]. Coastalwaters [Cla79]. code [Stu79]. COFI [Hju77]. College [Low78, Sha79a]. Colombo [Law77b]. COMECON [Kwa77, Hea77]. Commission [BS79, Gam77, Sco78]. commissions [Cus77]. Committee [Hju77]. Communications [Cus79]. community [You77]. company [Ion77]. Comparative [O’C78, Wat79a]. compared [Arc77]. compensation [Lof78b]. Competent [KM79]. compiled [O’C78]. Concepts [Gon79, Jos77]. conceptual [Kel78]. conduct [Stu79]. Conference [Kim77a, Pri77b, Sta77, Gil79, Kim77b, Kim79a, Low78, Sha79a, Wat77b]. conferences [Stu79]. conflict [MP78, Mit77]. Conflicts [Pul79b]. Congress [Pul77a]. Consent [Fri77a]. Conservation [Ano77a, Pri77b, Gam77, Kel78, Kni78, Tet78]. considerations [Sta78]. Constraints [Nas79]. construction [Sha79b]. Consultants [Met79]. continental [Len79, Wat79d]. continues [JG78]. control [Car78a, Pri77a, Seb79]. convenience [Abr79, Low77a]. convention [Hor77]. Cooperation [Bir77a, Pul79b]. cooperative [Ram77]. Coordinating [Sal78]. Council [Nas79, Ano78c]. countries [Ano78c, Kwa77, She78, Arc78]. crisis [Ric79c]. Croom [O’C77, You77]. cross [Gau79]. cross-boundary [Gau79].

D [Cas77, Por77]. Dahl [Chu78b]. Danish [Arc77]. Danny [Arc78]. data [AH79]. David [Pul79a]. DC [Abe77, Bir78a, Nas79]. debate [JG78]. Decade [Lof77a]. December [Law77b, Pul77a, Sta77]. decentralized [Kin79]. decision [Kin79]. deeper [Sta77, Wat77b]. Deepsea [Fri77b]. delimitation [Jac79]. demand [Gil79]. Department [Rus78]. design [Sha79b]. detailed [Chu77a]. Developing [Ano78a, Arc78, Law77a, She78]. Development [Law77b, Ram77, Col77b, Ion77, Jon78, McK79, Sim78, Stu79, Wat77b]. Df11 [Daw78]. Dfl [Chu77b]. directions [Law77a, O’C78]. disarmament [Lar79]. discontent [Lof78c]. dispute [Jac79]. disputes [Low77b]. distributed [Bir77a, Wat77a]. divergent [Ion77]. divers [Lof79e]. Diving [Mat78]. Division [Cus79]. do [Lof79c]. Doganis [Low77a]. door [Lof78b]. Douglas [You77]. down [Kim79b]. Drewry [Met79]. dritten [Wat79c].

East [Met79, Buz79, TL78]. Ebenhausen [Wat79c]. economic [Arc78, Chu77a, Chu78a, Dar77, Kin79, Met79, Ogl77, Sto79, Mei78]. economics [Ano78a, Gil79, Gla79]. ecosystem [Tet78]. ed [Rus78]. edge [Pul77b]. Edinburgh [Bir77b, Yon77]. edited [Chu78b, Len77b, Len77a, Len78, Mei78, Mor78, O’C78, O’C79, Pri77a, Rus78, Wat78, You77]. edition [Len78, Pul77b]. editor [Hea77, JG78]. Editorial [Ano77b, Ano78b, Ano79e]. Edward [Por77]. Edwards [Pul79a]. EEC [Chu77c, Pan79, Rey77, Riz78]. effort [Sto79]. Eggenberg [Wat79c]. Eighth [Kim79a]. Einar [Chu78b]. Electrical [Abe77]. Electronics [Abe77]. Eleventh [Hju77]. Elisabeth [Ogl77, Ogl79]. Elsevier [Daw78]. emerging [CS79]. enclosure [Wat79d]. Endel [You78b]. Endel-Jakob [You78b]. energy [Sha77a, Sha77b]. Enforcement [Car78a, Bir77a]. Engineering [Rus78, Daw78]. Engineers [Abe77]. Englewood [Sha79b]. enterprise [Law77b, You78a]. enterprises [Ogl79]. Environment [AR77]. Environmental [Bla77, Fre77, Rus78]. equipment [Lof78d]. essential [RV77]. Establishing [Bir78a]. establishment [Ano77a]. Ethel [Len77b]. Europe [Pul77a, Pul77c, Sha79a, Len79]. European [Len77b, Pul77a, Pul79b]. Exhibition [Sta77, Lof78d]. experience [Tom79]. exploitation [Lof78d, RS79, Gau79]. exploration [Bla77]. Extended [Lof79a, Mac79, Mei78]. extension [Col77b].

factors [Ber77, Wat78]. failure [Ell79]. Famborough [Wat79a]. Farmer [Pul77a]. Farming [Pul77a, Pul79a, Pul77b]. Farnborough [Wat77a]. Farnham [Pul77b, Pul79a]. feasibility [Chu77b]. federal [Low77b, McK79, Met78]. Ferenz [Wat79c]. field [Gau79]. fields [Ion77]. Fifth [Kim77a]. Fighting [Sie77]. Financial [Sta77]. Financing [Man79]. First [Wat79d, Lof79e]. Fish [Pul77a, Col77a, Gul77a, Hju77, Hol78, TL78]. Fisheries [Col77b, Hju77, Ric79a, Bir78a, Bir78b, CS79, Chu77c, Cus77, Gul77a, Law77a, Lof79a, Mei78, Par79, Pul79b, Sim78, Cus79]. Fishery [Chu78a, Car78a, Gla79, Kel78, Kni78, Law77b, WMCS77]. Fishing [Pul77b, Chu77a, Jos78, Jos79, Mac79, MP78, Ric79b, Sto79, Pul79a]. Fiskerigrensen [Chu78a]. flags [Low77a, Abr79]. fleets [Riz78]. Fleischer [Chu78a]. Floating [Kie77, Woo78]. floor [Wat77b]. Florida [Kau79]. Folkeretteim [Chu78a]. Food [Hju77, Law77b]. foreign [O’C77, Ric79b]. Forlag [Chu78b]. Forschungsinstitut [Wat79c]. Forum [Low78, Abe77, Sha79a]. framework [Kel78]. France [Lof78d, Sco78]. Francis [Wat79b]. Frederick [She78]. French [dC77]. Frigg [Gau79]. fund [Lof78b]. further [Woo78]. future [Ano78s, Met79, Par79].

G [Cas77, Mei78]. gains [Kim78]. Gamble [Wat79a]. gas [Ano77p, Gau79, Len77b, Len79]. Gaynanov [Cas77]. General [Ano78c]. Geneva [Kim79a]. given [Ogl79]. Goldberg [Por77]. Government [Bir78a, Ion77]. Grace [Sha79b]. Graham [Pri77a]. Grant [Bir77a, Cus79]. Greece [Pan79]. Greek [Pan79]. Greenwich [Low78, Sha79a, Sha79a].

Group [Chu77b, Low77a]. growing [Ano77o]. Grundey [Abr79]. Guildford [Wat78]. Gulf [Ano77a].

H [Met79]. Hall [Sha79b]. Handling [Col77a]. Hans [Wat79c]. Hants [Wat77a, Wat79a]. harbours [Lof78d]. harvest [Hju77]. Havretts [Chu78b]. Hawaii [Bar78]. Heikoff [Cla78]. Helm [O’C77, You77]. Henry [Cus79, Met78]. herring [Lof77c]. highly [Jos77, Kin79]. Hodder [O’C79]. Hole [Gul77b]. Holst [Wat78]. Honolulu [Bar78]. Houston [O’C78]. Hundred [Jos78]. Hydrographic [Has77, RV77]. Hydronamic [Chu77b]. Hypothetical [Bro78a]. Hysvaer [Chu78b].

ICES [Par79, TL78]. ideal [Sha77b]. impact [Buz79, Low77a, You77]. Impacts [RS79, Hea77, Mei78]. implications [Hea77, Hju77]. important [Jos77]. Improved [AH79]. improvements [RV77]. including [Ano77a, Chu77b]. independence [Kei77]. Index [Ano77c, Ano78d, Ano79f]. Indian [Ano77a]. Industrial [Chu77b]. industry [Jos78, Jos79]. inevitable [Fri77a]. Information [Har77, Fre77]. inland [Pri77a]. Inshore [Mac79]. Institute [Abe77, Bar78, Bir77a, Col77a, O’C78, Ogl77, Ogl79, Car78b]. Institutions [Cla78]. Insurance [Man78]. interdisciplinary [Fro77]. interest [Her78]. interests [Arc78, Ion77, Mac79]. Intergovernmental [Sco78]. International [Ano77a, BS79, Kin79, Low77b, Mas77, O’C78, Ogl79, Pri77b, Pul77a, Sha79b, Sta77, Abe77, Bir77a, Car78a, CS79, Chu78a, Gau79, Hea77, Hol78, KM79, Lof78d, Ogl77, Sha79a, Wat79b, Gam77]. internationalen [Wat79c]. introduction [Ano78c]. investment [Jos78, Jos79]. IPC [Wat78]. Iran [Ano77a]. island [And78, Chu77b, And78]. island-studded [And78]. Islands [Jac79]. Isles [Ric79a]. Israel [Car77, Car78b]. issues [Ano78s, Buz79, Lof79b]. Italy [Hju77]. Ivar [Chu78b]. Iversen [Pul77b]. IWC [Ell79].

J [Pri77a, Pul79a, Wat78, Wat79b, You78a]. Jakob [You78b]. James [Arc78, Met78]. Jane [Len77a, Len78, Len77a, Len78]. Janis [Wat77a]. January [Ano77i, Ano78k, Ano79h, Ano79f]. Japan [Pri77b, Sak79]. Japanese [Jos78, Jos79]. Jeremiah [You78a]. Jersey [Sha79b]. Johan [Wat78]. John [Cla79, Met78, Tet78, Wat79a, Por77, Tet78]. Johnston [Mor78, You77]. Joint [Yon77]. JONSDAP [Ram77]. Joseph [Cla78]. Jr [Wat79a]. July [Ano77h, Ano78j, Ano79f, Ano79g, Col77a, Kim77b]. June [BS79, Gil79, Kau79, Wat77b]. jurisdiction [Bro78b, Bro78c, Mac79, Par79, Mei78].

Karl [Chu78b]. Ken [Abr79]. Kenneth [Ion77]. Key [Kau79]. Keyser [Len77b]. King [Wat79a]. knowledge [Fre77]. Kogan [Len77b, Len79, Ion77]. Kolde [You78b]. krill [Lof78a]. Kuypers [Met78].

L [Arc78, Len77a, Len78, Rus78]. lands [Low77b]. Lanka [Law77b]. Latin [Pal79]. Law [Bar78, Kim77a, Lof79b, Lof79c, And78, Bir77a, CS79, Chu78a, Chu78b, Fri78, Gon79, Jac79, Kie77, Kim77b, Kim79a, KM79, Lar79, Low77b, O’C78, Ogl77, Rey77, Wat77a, You77, You78a, You79, Jos78]. Lay [O’C78]. leads [Lof77b]. Lee [Mei78]. legal [Sta78, Woo78]. Legatski [Bir77a]. legislation [Fri77b, Mat78]. Leipziger [Arc78]. Letters [Hea77, JG78]. level [Wat79b]. Lexington [Wat77a, Wat79a, You78b, You78a]. Liability [Fit78]. light [Par79]. Limitation [Sto79]. Limited [Pul79a]. Limits [Chu78a, Bro78b, Bro78c, Chu77a, Lof79a]. liner [Met79, Stu79]. Living [Gul77b]. LNG [Man78, Man79]. London [Abr79, Bir77a, BS79, Col77a, Ion77, Len77b, Len77a, Len78, Len79, Low77a, Low78, Met79, Mor78, O’C77, O’C78, O’C79, Por77, Pri77a, Pul77a, Tet78, Wat77b, Wat79b, You77]. Ltd [Len79, Met79, Wat77a].

M [Arc78, Cas77, Cla78, Cla79, O’C78, You77]. MA [Wat79a, You78b, You78a]. MacDonald [Len78, Len77a]. making [Kin79]. Malta [Ogl79, Ogl77]. Management [Bir78b, Kni78, Cla78, Cla79, Fre77, Gul78, Jos77, JG78, Kel78, Kin79, Law77b, RS79, Sak79, Tet78, Wat79b, WMCS77]. Managing [Hol78]. manganese [AR77]. Mann [Ogl77, Ogl79]. manual [Tet78]. March [Ano77a, Kim79a]. Marcus [Met78]. Mariculture [Pul77a]. Marine [Abe77, Ano77a, Bir77a, Car77, Car78b, Cus79, Jon78, Pri77b, Sno78, Fri77a, Gon79, Guz77, Hor77, Por77, Val78, Mor78, Sha79b, Wat79a]. Maritime [Buz79, Ano78a, You77]. Mark [Wat77a]. market [Ano77o]. Mass [Arc78, Gul77b, Met78, She78]. May [Bir77b, Kau77, Kau79, Kim77b, Pri77b, Sha79a]. medical [Kau77, Kau79]. Mediterranean [Hol78, Sal78]. meeting [Bar78, BS79, Kau77, Kau79, Bir77b]. Meetings [Ano77d, Ano77e, Ano77f, Ano77g, Ano78e, Ano78f, Ano78g, Ano78h]. member [Ano78c, Riz78]. merchant [Guz77, Riz78]. Merges [Pri77b]. Metaxas [Low77a]. MI [Cla79]. Mich [Cla78]. Michigan [Mei78]. Middle [Met79]. migratory [Jos77, Kin79]. mile [Bir78a, Chu77a, Jos78]. Military [You79]. mineral [Arc78]. mining [AR77, Fri77b, Lof78e, Mar79]. MIR [Cas77]. Molehill [Lof78e]. momentum [Kim78]. Monitoring [TL78, Por77]. Morges [Ano77a]. Moscow [Cas77]. mountain [Lof78e]. Moving [Wat77b]. Mudge [Arc78]. München [Wat79c].

N [Low77a]. Nandrup [Chu78b]. National [Nas79, Pul77a, Bro78b, Bro78c, Bir77a]. Nationen [Wat79c]. Nations [Hju77, Kim77a, Kim77b, Kim79a]. Natural [Ano77a, Pri77b]. Nature [Ano77a, Pri77b]. Naval [Low78, Sha79a, O’C79]. Navies [O’C77]. neglected [Lof79b]. Nes [Chu78b]. Nesmeyanov [Cas77]. News [Pul77b, Pul79a]. NKr69 [Chu78a]. No [Ogl77, Ogl79, Pre78, You78b, You78a]. nodule [AR77]. Nodules [Mar79]. Nord [Wat79c]. Nordquist [O’C78]. Norges [Chu78b]. Norsk [Chu78b].

North [Chu77b, Ion77, Wat78, Wat79b, Arc77, Ber77, Bla77, Bro78a, Buz79, Fit78, Lof77b, MP78, Mas77, Pri78, Ram77, Sha79a, TL78]. North-East [Buz79, TL78]. Northern [Ano77a]. Norway [Chu78b, Pul79a]. Norwegian [Arc77]. November [Bar78, Law77b, Ogl79, Pul77a, Sco78]. Nr [Wat79c].

O [Met78]. occasion [Gil79]. Occasional [Ogl77, Ogl79]. Ocean [Ano77a, Fre77, Ogl77, Ogl79, AH79, Daw78, Len77a, Len78, Lof78e, Gul77b, Bir77a]. Oceanographic [Sco78, Yon77]. oceanography [Daw78]. Oceans [O’C78, Kei77, Lof77a, Lof78d, McK79, Wat79d, Cas77]. Oceanus [Pre78]. October [Ano77j, Ano78l, Ano79f, Ano79i, Lof78d, Low78, Sco78, Wat79c]. Odell [Ion77]. Office [Bir78a]. Offshore [Ano78i, Raw78, Sta78, Sta77, Ano77o, Len77b, Lof78d, Low77b, Len79]. Oil [MP78, Ano77p, Bro78a, Fit78, Ion77, Len77b, Len79, Lof77b, Pri77a, Sta77]. økonomiske [Chu78a]. Ontario [Kau77]. operational [Sta78]. operations [Bla77]. opportunities [Mar79, Nas79, Kei77]. Optimal [Ion77]. options [Sto79]. Oran [Wat79b]. order [Ogl77]. Organisation [Hju77, Law77b]. organised [Sta77]. organization [Gul78]. Organizational [Kel78]. organizations [Hol78, KM79]. organized [Gil79, Wat77b]. origins [Ram77, Wat79d]. Oslo [Chu78a, Chu78b, Wat78]. Ostsee [Wat79c]. outfall [Sha79b]. overseas [Abr79]. Oxford [Daw78].

P [Met79]. Pacific [You78b, You78a, Wat79b, You78a]. Page [Ion77, Len77b, Len79]. pages [Ano77a, Arc78, Bir77a, Cas77, Chu78a, Daw78, Ion77, Len77b, Len77a, Len78, Low77a, Mei78, Met78, Mor78, O’C77, Ogl77, Por77, Pri77a, Pri77b, Pul77b, Rus78, Wat77a, You77, Ano77k, Ano77i, Ano77h, Ano77j, Ano78m, Ano78k, Ano78j, Ano78l, Ano79j, Ano79h, Ano79g, Ano79i, Bir78a, Chu78b, You78b, You78a]. Paper [Ogl77]. Papers [Ogl79, Ano78r]. parcel [Lof77d]. Pardo [Ogl77]. Paris [Sco78]. Parks [Ano77a, Pri77b]. Parliament [Seb79]. Part [Bro78b, Bro78c]. particular [Cus79]. patrol [Ano78i]. pattern [Jon78]. Paul [Met78]. Persian [Ano77a]. perspective [dC77]. Peruvian [Gla79]. Peter [Ion77, Cla79]. petroleum [Gau79, Cas77]. philosophy [Ram77]. Pinter [Wat79b]. Pity [Lof77c]. Plains [Sno78]. Planning [Sha79b, Ram77]. Platzöder [Wat79c]. policeman [Raw78]. policemen [Sta78]. Policies [Sak79, Arc77, Pal79]. policy [Chu77c, Chu78b, Fre77, Fri78, Hea77, Kwa77, Met78, O’C79, Ric79a, Ric79b, Sha77a, Wat79a, You77, O’C77]. Politics [WMCS77, Buz79, Fri78, Gla79, Mas77]. Politik [Wat79c]. politische [Wat79c]. Pollution [dC77, Dar77, Fit78, Hor77, Por77, Pri77a, Sak79, Seb79, Sie77, Mor78]. Polytechnic [Abr79, Low77a]. poor [Lof77c]. port [Met78]. ports [Met79]. postage [Chu77b]. Potential [Mit77]. power [Wat77a]. pp [Abr79, Chu77b, Cla78, Cla79, Cus79, Len79, Ogl79, Pul79a, Sha79b, She78, Tet78, Wat78]. pp. [Met79, Wat79a, Wat79b, Wat79c]. practice [Fre77]. Prentice [Sha79b]. Prentice-Hall [Sha79b]. Press [Bir77a, Met78, Mor78, Ogl79, Wat78]. price [Ogl79]. Printing [Bir78a]. problems [Fro77, Hju77, Car78a, Hor77]. procedures [Kni78]. proclamation [Wat79d]. production [Bla77]. Products [Col77a]. Professor [Gil79]. programs [Cla78]. proliferation [Chu77a]. Promotion [Ano77a]. Prospects [Val78, Met79]. Protecting [Sal78]. protection [Bla77]. public [Her78]. Publications [Ano77l, Ano77m, Ano77n, Ano78n, Ano78o, Ano78p, Ano78q, Ano79k, Ano79l, Ano79m, Ano79n, Met79, Wat79c]. published [Bir77a]. Publishers [Cas77, Cla79].

quest [You78b].

R [Ion77, Low77a, O’C78]. R. [Mor78]. Radioactive [Pre78]. Ranft [O’C79]. Ray [She78]. reconciliation [Ion77]. Red [Ano77a]. reference [Cus79]. Regier [Cus79]. regime [Fri77a, Sha79a]. region [Gon79]. regional [Buz79, Sal78]. regionalism [Val78]. regulations [Car78a]. reinsurance [Man78]. relation [Jos78]. Renate [Wat79c]. renewable [Cus79]. Report [Bir78a, Chu77b, Low77a]. requirements [Man78]. Research [Low77a, Nas79, You78b, You78a, Car77, Dar77, Fri77a, Ram77, Sim78]. reserve [Lof79d]. reserves [Ano77a, Pri77b]. resource [RS79, Wat79b]. Resources [Ano77a, Cus79, Dar77, Lof78c, Lof79d, Mit77, Pri77b, Arc78, Cas77, Chu78b, Cla78, Gul77b, Wat79d, Tet78]. resourceswith [Cus79]. response [Mac79]. ressurspolitikk [Chu78b]. retirement [Gil79]. Review [Abr79, Ano77a, Arc78, Bir77a, Bir78a, Cas77, Chu77b, Chu78a, Chu78b, Cla78, Cla79, Cus79, Daw78, Ion77, Len77b, Len77a, Len78, Len79, Low77a, Mei78, Met78, Met79, Mor78, Nas79, O’C77, O’C78, O’C79, Ogl77, Ogl79, Por77, Pri77a, Pri77b, Pul77b, Pul79a, Rus78, Sha79b, She78, Tet78, Wat77a, Wat78, Wat79a, Wat79b, Wat79c, You77, You78b, You78a]. revision [Chu77c]. Richard [Bir77a]. Rim [You78b, You78a]. Robert [Len77a, Len78, Sha79b]. Roberts [Met78]. Rockall [Bro78b, Bro78c]. role [Col77b, Hol78, McK79, Par79, Sim78, TL78, Gam77]. Rome [Hju77]. Rosing [Ion77]. Royal [Low78, Wat77b, Sha79a]. rural [Col77b]. Ryner [Cla79].

S [Low77a, Met78, O’C78, Pul77b, Wat79c]. Safety [Lof79e]. Salmon [Pul79a]. scale [Law77a, Law77b, Lof79a]. Scandinavia [Mar79]. SCAR [Gul77b]. school [Gil79]. Science [Cla78, Cla79, Gla79, Mei78, Wat78, Cus79]. Sciences [Nas79]. Scotland [Bro78a]. Sea [Ano77a, Bar78, Bir77a, Chu77b, Cus79, Ion77, Kim77a, And78, Bir78b, Chu78b, Fri78, Gon79, Has77, Jac79, Kim77b, Kim79a, KM79, Lar79, Lof77b, Lof78d, Lof79b, Lof79c, Pri77a, Rey77, Sha79a, Wat77b, You79, Arc77, Bla77, Bro78a, Fit78, MP78, Mas77, Pri78, Ram77, Sha79a, Val78, O’C78, Ogl77, Pul77b, Wat77a, You77, You78a, Wat77a]. seabed [Pre78, Arc78]. seas [Cas77]. Seattle [Bir77a, Cus79]. Second [Pul77a, Sta77, Pul77b]. Security [Lar79]. Seerechtskon [Wat79c]. self [Ano77p]. self-sufficiency [Ano77p]. September [Abe77, Kim77a, Wat79d, Yon77]. Seregin [Cas77]. Series [You78b, You78a]. service [Har77]. services [AH79, Col77b, Has77, Met79, RV77]. Session [Hju77, Kim77a, Kim77b, Kim79a, Sco78]. settlement [Jac79]. Severn [Rus78]. Shaw [Rus78]. shelf [Wat79d]. shipbuilding [Lof78d]. Shipping [Kwa77, Guz77, Hea77, Met79, Pan79]. shipyards [Ric79c]. Short [Met78]. should [Lof79c]. Sicherheit [Wat79c]. situation [Ano78c]. Sixth [Kim77b]. Sliedrecht [Chu77b]. small [Law77a, Law77b, Lof79a]. small-scale [Law77a, Law77b, Lof79a]. Smith [Pri77a]. Society [Abe77, Pul77a, Wat77b, Kau77, Kau79]. Sokolov [Cas77]. solutions [Hor77]. Some [Jos77]. something [Lof79d]. sone [Chu78a]. Sons [Por77, Tet78]. source [Sha77b]. South [Dar77, Val78]. Southern [Gul77b]. Soviet [Guz77]. space [And78]. Spain [Ric79b]. Spanish [Ric79c]. special [Pri78]. species [Jos77, Kin79]. sponsored [Abe77, Law77b, Pul77a]. Sri [Law77b]. stable [Lof78b]. states [And78, Riz78, Met78, Nas79]. status [Woo78]. steps [Wat79d]. Stiftung [Wat79c]. Stockholm [Gil79]. stocks [Gul77a, Hol78, TL78]. Stoughton [O’C79]. straits [And78]. strategic [Ber77, Wat78]. Strategies [Por77]. Street [Met79]. structure [Kel78]. studded [And78]. Studies [Bir77a, Low77a]. study [Bro78a, Ion77, Chu77b]. Submersibles [Daw78]. sufficiency [Ano77p]. Sullivan [You78a]. Summaries [Ano78r, Ano78c]. supply [Gil79]. Surrey [Pul77b, Pul79a]. survey [Chu77a]. Sweden [Gil79]. Switzerland [Kim79a, Ano77a, Pri77b]. SWP [Wat79c]. SWP-S [Wat79c]. Sydney [Por77]. Symposium [Gul77b]. systems [Sha79b].

T [Bir77a, Rus78]. tabulated [Ano78c]. target [Ano77p]. Teakfield [Wat77a]. technical [Ano78r, Tet78, O’C79]. techniques [Lof78d]. Technology [Abe77, Bir78a, Wat78, Len77a, Len78]. Tehran [Ano77a]. Tenth [Sco78]. their [Daw78, Ion77, Woo78, Buz79, Col77b, Mac79]. theory [Fre77]. Third [Lof78d, Kim77b, Len78, Low78]. Thomas [Gil79]. Thorburn [Gil79]. thoughts [Woo78]. threat [Lof79a]. Three [Jac79]. Tidal [Sha77b, Sha77a]. Tiden [Chu78b]. Times [Sta77]. Tokyo [Pri77b]. Topics [Ano78s]. Toronto [Por77, Kau77]. traffic [Met79]. transnational [Kie77]. Transport [Low77a, Gil79]. treaty [Riz78]. Treholt [Chu78b]. triennial [Lof78d]. Trillo [Len77a, Len78]. tropical [Col77a, Col77a]. Trotman [Pri77a]. trout [Pul79a]. Truman [Wat79d]. tuna [Gul78, JG78]. Two [Arc77, Jos78]. Two-Hundred-Mile [Jos78].

UK [Abr79, Bir77b, Col77a, Len79, Low78, Pul77a, Pul79a, Sta77, Wat77a, Wat77b, Yon77, Ano77p, Ano77o, BS79, Bro78b, Bro78c, Har77, Jon78, Len77a, Mat78, Seb79, Sha79a, You77, Low77a]. UNCLOS [Kim78, Kim79b, Lof77d, Lof78e]. Undersea [Kau77, Kau79]. Union [Ano77a, Pri77b, Pul77a]. United [Hju77, Kim77a, Kim77b, Met78, Nas79, Kim79a]. Universitetsforlaget [Chu78a, Wat78]. University [Bir77a, Cus79, Ogl79, Rus78]. unpacked [Lof77d]. update [Jos79]. USA [Cla79, Gul77b, Kau79, Kim77a, Kim77b, Sha79b, Kei77]. use [Daw78]. users [Has77].

V [Cas77, Seb79]. Vereinten [Wat79c]. versus [Kin79]. vessels [Ano78i]. vital [Has77]. Vol [Pre78]. Volkerrechtliche [Wat79c]. Vols [O’C78]. volume [Ano77c, Ano78d, Ano79f, O’C78].

W [She78, Wat77a]. W1Y2LL [Met79]. WA [Cus79]. Wardley [Pri77a]. Wardley-Smith [Pri77a]. Washington [Abe77, Bir77a, Bir78a, Cus79, Nas79, Cus79]. wastes [Pre78]. water [Gil79]. waters [Sta77, Wat77b, Pri77a]. Whale [Gam77]. Whaling [BS79, Gam77, Bir77b]. Whitehall [Seb79]. Who [Lof77b, Lof79c]. whom [Lof77b]. Wiley [Por77, Tet78]. William [Bir77a]. wire [Kim79b]. Wissenschaft [Wat79c]. Woodhead [Bir77a]. Woods [Gul77b]. Workshop [Law77b]. World [Gul77a, Hju77, Guz77].

yearbook [Len77b, Len79]. York [Daw78, Kim77a, Kim77b, Mor78, Por77, Tet78, Wat79b]. Young [Wat79b].

zone [Chu78a, Dar77, Kni78, Tet78, Bir78a]. zones [Chu77a].

References

Abel:1977:IFS

[Abe77] Robert B. Abel. An international forum sponsored by the Marine Technology Society and the Institute of Electrical and Electronics Engineers: Washington, DC, 13–15 September 1976. Marine Policy, 1(2):165–168, April 1977. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

          com/science/article/pii/0308597X77900525.

Abrahamsson:1979:BRF

[Abr79] B. J. Abrahamsson. Book review: Flags of convenience in 1978: Ken Grundey, Polytechnic of Central London, London, 1978, 58 pp, £3.00 (UK), £5.00 (overseas). Marine Policy, 3 (4):327, October 1979. CODEN ???? ISSN 0308-597X (print),

          1872-9460 (electronic). URL http://www.sciencedirect.com/
          science/article/pii/0308597X79900290.

Abram:1979:IAO

[AH79] Richard J. Abram and Kent H. Hughes. Improved access to ocean data services. Marine Policy, 3(4):315–317, October 1979. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X7990023X.

Andrew:1978:ALS

[And78] Dale Andrew. Archipelagos and the law of the sea: Island straits states or island-studded sea space? Marine Policy, 2(1): 46–64, January 1978. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0308597X7890060X.

Anonymous:1977:BRP

[Ano77a] Anonymous. Book review: Promotion of the establishment of Marine Parks and reserves in the Northern Indian Ocean including the Red Sea and Persian Gulf (Tehran, Iran, March 1975): 169 pages, $8.50, 1976, International Union for Conservation of Nature and Natural Resources, Morges, Switzerland. Marine Policy, 1(1):86, January 1977. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0308597X77900203.

Anonymous:1977:EB

[Ano77b] Anonymous. Editorial Board. Marine Policy, 1(1):ifc, January 1977. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X7790001X.

Anonymous:1977:IV

[Ano77c] Anonymous. Index to volume 1. Marine Policy, 1(4):363– 364, October 1977. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0308597X77901051.

Anonymous:1977:Ma

[Ano77d] Anonymous. Meetings. Marine Policy, 1(1):88, January 1977. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0308597X77900239.

Anonymous:1977:Mb

[Ano77e] Anonymous. Meetings. Marine Policy, 1(2):176, April 1977. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0308597X77900598.

Anonymous:1977:Mc

[Ano77f] Anonymous. Meetings. Marine Policy, 1(3):271–272, July 1977. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X77900422.

Anonymous:1977:Md

[Ano77g] Anonymous. Meetings. Marine Policy, 1(4):361, October 1977. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X7790104X.

Anonymous:1977:PJb

[Ano77h] Anonymous. Pages 178-272 (July 1977). Marine Policy, 1(3):??, July 1977. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:1977:PJa

[Ano77i] Anonymous. Pages 2–88 (January 1977). Marine Policy, 1(1):??, January 1977. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:1977:PO

[Ano77j] Anonymous. Pages 274-364 (October 1977). Marine Policy, 1(4):??, October 1977. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:1977:PA

[Ano77k] Anonymous. Pages 90-176 (April 1977). Marine Policy, 1(2):??, April 1977. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:1977:Pb

[Ano77l] Anonymous. Publications. Marine Policy, 1(2):174, April 1977. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X77900574.

Anonymous:1977:Pc

[Ano77m] Anonymous. Publications. Marine Policy, 1(3):270, July 1977. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0308597X77900409.

Anonymous:1977:Pd

[Ano77n] Anonymous. Publications. Marine Policy, 1(4):357–358, October 1977. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X77901026.

Anonymous:1977:UOM

[Ano77o] Anonymous. UK offshore market: a growing adolescent. Marine Policy, 1(1):75–76, January 1977. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X77900136.

Anonymous:1977:UOG

[Ano77p] Anonymous. UK oil and gas: on target for self-sufficiency. Marine Policy, 1(3):258–259, July 1977. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X7790032X.

Anonymous:1978:DME

[Ano78a] Anonymous. Developing maritime economics. Marine Policy, 2(4): 337–339, October 1978. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0308597X78900258.

Anonymous:1978:EB

[Ano78b] Anonymous. Editorial Board. Marine Policy, 2(1):ifc, January 1978. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X78900544.

Anonymous:1978:GIT

[Ano78c] Anonymous. General introduction and tabulated summaries of the situation in member countries of the Council. Marine Policy, 2(2):112–126, April 1978. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X78900040.

Anonymous:1978:IV

[Ano78d] Anonymous. Index to volume 2. Marine Policy, 2(4):347– 348, October 1978. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0308597X78900350.

Anonymous:1978:Ma

[Ano78e] Anonymous. Meetings. Marine Policy, 2(1):83–84, January 1978. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X78900726.

Anonymous:1978:Mb

[Ano78f] Anonymous. Meetings. Marine Policy, 2(2):ibc, April 1978. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0308597X78900155.

Anonymous:1978:Mc

[Ano78g] Anonymous. Meetings. Marine Policy, 2(3):250, July 1978. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0308597X78900532.

Anonymous:1978:Md

[Ano78h] Anonymous. Meetings. Marine Policy, 2(4):344–345, October 1978. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X78900337.

Anonymous:1978:OPV

[Ano78i] Anonymous. Offshore patrol vessels. Marine Policy, 2(3):237– 239, July 1978. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0308597X7890043X.

Anonymous:1978:PJb

[Ano78j] Anonymous. Pages 170-250 (July 1978). Marine Policy, 2(3):??, July 1978. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:1978:PJa

[Ano78k] Anonymous. Pages 2–84 (January 1978). Marine Policy, 2(1):??, January 1978. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:1978:PO

[Ano78l] Anonymous. Pages 254-348 (October 1978). Marine Policy, 2(4):??, October 1978. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:1978:PA

[Ano78m] Anonymous. Pages 86-168 (April 1978). Marine Policy, 2(2):??, April 1978. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:1978:Pb

[Ano78n] Anonymous. Publications. Marine Policy, 2(1):81–83, January 1978. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X78900714.

Anonymous:1978:Pc

[Ano78o] Anonymous. Publications. Marine Policy, 2(2):167–168, April 1978. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X78900143.

Anonymous:1978:Pd

[Ano78p] Anonymous. Publications. Marine Policy, 2(3):240–241, July 1978. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X78900453.

Anonymous:1978:Pe

[Ano78q] Anonymous. Publications. Marine Policy, 2(4):342–343, October 1978. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X78900313.

Anonymous:1978:STP

[Ano78r] Anonymous. Summaries of technical papers. Marine Policy, 2 (2):151–157, April 1978. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0308597X78900088.

Anonymous:1978:TFI

[Ano78s] Anonymous. Topics for future issues …. Marine Policy, 2(4): 346, October 1978. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0308597X78900349.

Anonymous:1979:Ca

[Ano79a] Anonymous. Calendar. Marine Policy, 3(1):75, January 1979. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0308597X79900484.

Anonymous:1979:Cb

[Ano79b] Anonymous. Calendar. Marine Policy, 3(2):167, April 1979. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0308597X79900149.

Anonymous:1979:Cc

[Ano79c] Anonymous. Calendar. Marine Policy, 3(3):252, July 1979. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0308597X79900654.

Anonymous:1979:Cd

[Ano79d] Anonymous. Calendar. Marine Policy, 3(4):ibc, October 1979. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0308597X79900344.

Anonymous:1979:EB

[Ano79e] Anonymous. Editorial Board. Marine Policy, 3(1):ifc, January 1979. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X79900356.

Anonymous:1979:IVJ

[Ano79f] Anonymous. Index to volume 3: January, April, July, and October 1979. Marine Policy, 3(4):331–332, October 1979. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X79900332.

Anonymous:1979:PJb

[Ano79g] Anonymous. Pages 170-252 (July 1979). Marine Policy, 3(3):??, July 1979. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:1979:PJa

[Ano79h] Anonymous. Pages 2–75 (January 1979). Marine Policy, 3(1):??, January 1979. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:1979:PO

[Ano79i] Anonymous. Pages 254-332 (October 1979). Marine Policy, 3(4):??, October 1979. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:1979:PA

[Ano79j] Anonymous. Pages 78-167 (April 1979). Marine Policy, 3(2):??, April 1979. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:1979:Pb

[Ano79k] Anonymous. Publications. Marine Policy, 3(1):74–75, January 1979. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X79900472.

Anonymous:1979:Pc

[Ano79l] Anonymous. Publications. Marine Policy, 3(2):165–166, April 1979. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X79900137.

Anonymous:1979:Pd

[Ano79m] Anonymous. Publications. Marine Policy, 3(3):250–251, July 1979. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X79900630.

Anonymous:1979:Pe

[Ano79n] Anonymous. Publications. Marine Policy, 3(4):329–330, October 1979. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X79900320.

Amos:1977:EAM

[AR77] Anthony F. Amos and Oswald A. Roels. Environment aspects of manganese nodule mining. Marine Policy, 1(2):156– 163, April 1977. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0308597X77900501.

Archer:1977:TNS

[Arc77] Clive Archer. Two North Sea policies: Danish and Norwegian attitudes compared. Marine Policy, 1(4):289–300, October 1977. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X77900859.

Archer:1978:BRS

[Arc78] Alan Archer. Book review: Seabed mineral resources and the economic interests of developing countries: Danny M. Leipziger and James L. Mudge, Ballinger, Cambridge, Mass, 1977, 241 pages, £11.00, $18.15. Marine Policy, 2(2):163–165, April 1978. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X7890012X.

Bardach:1978:AML

[Bar78] John E. Bardach. Annual meeting of the Law of the Sea Institute: Honolulu, Hawaii, 14–17 November 1977. Marine Policy, 2 (2):159–161, April 1978. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0308597X78900106.

Bertram:1977:NSF

[Ber77] Christoph Bertram. New strategic factors in the North Atlantic. Marine Policy, 1(1):37–44, January 1977. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X77900069.

Birnie:1977:BRN

[Bir77a] Patricia Birnie. Book review: National and international law enforcement in the ocean: William T. Burke, Richard Legatski, and William Woodhead published by Washington Sea Grant in Cooperation with the Institute for Marine Studies, and distributed by the University of Washington Press, Seattle and London, 139 pages, $7.50. Marine Policy, 1(2):172–173, April 1977. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X77900562.

Birnie:1977:MWE

[Bir77b] Patricia Birnie. Meeting on whaling: Edinburgh, UK, 13–14 May 1977. Marine Policy, 1(4):348–350, October 1977. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X77900963.

Birnie:1978:BRE

[Bir78a] Patricia Birnie. Book review: Establishing a 200-mile fisheries zone: Office of Technology Assessment Report, US Government Printing Office, Washington. DC, 1977, 132 pages. Marine Policy, 2(1):77–78, January 1978. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X78900684.

Birnie:1978:MSF

[Bir78b] Patricia Birnie. Management of sea fisheries. Marine Policy, 2 (3):235–237, July 1978. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0308597X78900428.

Blaikley:1977:EPN

[Bla77] D. R. Blaikley. Environmental protection in North Sea exploration and production operations. Marine Policy, 1(2):143– 155, April 1977. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0308597X77900495.

Brown:1978:SOH

[Bro78a] E. D. Brown. It’s Scotland’s oil?: Hypothetical boundaries in the North Sea — a case study. Marine Policy, 2(1):3– 21, January 1978. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0308597X78900568.

Brown:1978:RLNa

[Bro78b] E. D. Brown. Rockall and the limits of national jurisdiction of the UK: Part 1. Marine Policy, 2(3):181–211, July 1978. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X78900386.

Brown:1978:RLNb

[Bro78c] E. D. Brown. Rockall and the limits of national jurisdiction of the UK: Part 2. Marine Policy, 2(4):275–303, October 1978. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X78900192.

Birnie:1979:MIW

[BS79] Patricia Birnie and Richard Sandbrook. 30th meeting of the International Whaling Commission: 26–30 June 1978, London, UK. Marine Policy, 3(1):69–72, January 1979. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X79900435.

Buzan:1979:MIN

[Buz79] Barry Buzan. Maritime issues in North-East Asia: Their impact on regional politics. Marine Policy, 3(3):190–200, July 1979. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X79900514.

Carr:1977:MRI

[Car77] Judy Carr. Marine research in Israel. Marine Policy, 1(4):343– 344, October 1977. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0308597X77900938.

Cardoso:1978:EFR

[Car78a] J. C. Esteves Cardoso. Enforcement of fishery regulations: Problems in coastal and international control. Marine Policy, 2(2): 142–150, April 1978. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0308597X78900076.

Carr:1978:MII

[Car78b] Judy Carr. Marine institute in Israel. Marine Policy, 2(1): 68, January 1978. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0308597X78900623.

Caston:1977:BRP

[Cas77] V. N. D. Caston. Book review: Petroleum resources of seas and oceans: B. A. Sokolov, A. G. Gaynanov, D. V. Nesmeyanov and A. M. Seregin, MIR Publishers, Moscow, 1976, 267 pages, £2.50. Marine Policy, 1(4):353–354, October 1977. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X77900999.

Churchill:1977:MLR

[Chu77a] Robin Churchill. 200-mile limits: recent claims: a detailed annotated survey of the recent proliferation of claims to 200mile fishing limits and economic zones. Marine Policy, 1(3): 255–258, July 1977. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0308597X77900318.

Churchill:1977:BRI

[Chu77b] Robin Churchill. Book review: Industrial island in the North Sea Report on the feasibility study, The North Sea Island Group: Hydronamic, Sliedrecht, 1976, 195 pp, Dfl 55 (including postage). Marine Policy, 1(2):170–171, April 1977. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X77900550.

Churchill:1977:EFP

[Chu77c] Robin Churchill. The EEC fisheries policy: Towards a revision. Marine Policy, 1(1):26–36, January 1977. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X77900057.

Churchill:1978:BRF

[Chu78a] Robin Churchill. Book review: Fiskerigrensen, Folkeretteim og den økonomiske sone, (Fishery Limits, international law and the economic zone): by Carl A. Fleischer Universitetsforlaget, Oslo, 1977, 220 pages, NKr69. Marine Policy, 2(3):247–248, July 1978. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X78900519.

Churchill:1978:BRN

[Chu78b] Robin Churchill. Book review: Norges Havretts — og ressurspolitikk (Norway’s law of the sea and resources policy): edited by Arne Treholt, Karl Nandrup Dahl, Einar Hysvaer, and Ivar Nes, Tiden Norsk Forlag, Oslo, 1976, 215 pages. Marine Policy, 2(3):247–248, July 1978. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X78900507.

Clark:1978:BRC

[Cla78] M. J. Clark. Book review: Coastal resources management: Institutions and programs: by Joseph M. Heikoff, Ann Arbor Science, Ann Arbor, Mich, 1977, 287 pp. £11.00, $18.15. Marine Policy, 2(3):242–244, July 1978. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X78900477.

Clark:1979:BRC

[Cla79] M. J. Clark. Book review: Coastalwaters: a management analysis: by John M. Armstrong and Peter C. Ryner, Ann Arbor Science Publishers, Inc. Ann Arbor, MI. USA, 1978, 240 pp. £12.60. Marine Policy, 3(2):161–163, April 1979. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X79900101.

Cole:1977:HTF

[Col77a] R. C. Cole. Handling tropical fish: Tropical Products Institute, London, UK, 5–9 July 1976. Marine Policy, 1(1):77– 79, January 1977. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0308597X77900148.

Cole:1977:FES

[Col77b] Robert C. Cole. Fisheries extension services: Their role in rural development. Marine Policy, 1(2):132–142, April 1977. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X77900483.

Carroz:1979:NIL

[CS79] J. E. Carroz and M. J. Savini. The new international law of fisheries emerging from bilateral agreements. Marine Policy, 3 (2):79–98, April 1979. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0308597X79900022.

Cushing:1977:AFC

[Cus77] D. H. Cushing. The Atlantic fisheries commissions. Marine Policy, 1(3):230–238, July 1977. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X7790029X.

Cushing:1979:BRB

[Cus79] D. H. Cushing. Book review: A balanced science of renewable resourceswith particular reference to fisheries: by Henry A. Regier, Washington Sea Grant Communications, Division of Marine Resources, 3716 Brooklyn Avenue, University of Washington, Seattle, WA 98195, 1978. 108 pp, $9.50. Marine Policy, 3 (3):245–246, July 1979. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0308597X79900599.

Darracott:1977:RPR

[Dar77] Ann Darracott. Resources, pollution and research in a South African economic zone. Marine Policy, 1(3):239–254, July 1977. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X77900306.

Dawson:1978:BRS

[Daw78] James Dawson. Book review: Submersibles and their use in oceanography and ocean engineering: Elsevier, Amsterdam, New York and Oxford, 1977, 383 pages, $49.00, Df11 20.00. Marine Policy, 2 (4):342, October 1978. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0308597X78900301.

deCastro:1977:PFP

[dC77] Gene de Castro. Pollution — a French perspective. Marine Policy, 1(4):342–343, October 1977. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X77900926.

Elliot:1979:FI

[Ell79] G. H. Elliot. The failure of the IWC, 1946–1966. Marine Policy, 3(2):149–155, April 1979. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X7990006X.

Fitzmaurice:1978:LNS

[Fit78] Victor E. Fitzmaurice. Liability for North Sea oil pollution. Marine Policy, 2(2):105–111, April 1978. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X78900039.

Freeman:1977:OEI

[Fre77] Robert R. Freeman. Ocean and environmental information: the theory, policy, and practice of knowledge management. Marine Policy, 1(3):215–229, July 1977. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X77900288.

Fricke:1977:CRI

[Fri77a] Peter Fricke. Consent regime inevitable for marine research? Marine Policy, 1(1):71–73, January 1977. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X77900112.

Friedman:1977:DML

[Fri77b] Alan G. Friedman. Deepsea mining legislation. Marine Policy, 1(4):341–342, October 1977. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X77900914.

Friedman:1978:ULS

[Fri78] Alan G. Friedman. US law of the sea policy: a bureaucratic politics analysis. Marine Policy, 2(4):304–320, October 1978. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X78900209.

Frosch:1977:API

[Fro77] Robert Frosch. Anticipating problems — an interdisciplinary approach. Marine Policy, 1(2):164–165, April 1977. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X77900513.

Gambell:1977:WCR

[Gam77] Ray Gambell. Whale conservation: Role of the International Whaling Commission. Marine Policy, 1(4):301–310, October 1977. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X77900860.

Gault:1979:FGF

[Gau79] Ian Gault. The Frigg gas field: Exploitation of an international cross-boundary petroleum field. Marine Policy, 3(4):302– 311, October 1979. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0308597X79900204.

Gilman:1979:CSD

[Gil79] Sidney Gilman. The conference on supply and demand of water transport, organized by the Stockholm school of economics on the occasion of the retirement of Professor Thomas Thorburn: Stockholm, Sweden, 18–19 June 1979. Marine Policy, 3(4):322– 324, October 1979. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0308597X79900265.

Glantz:1979:SPE

[Gla79] Michael H. Glantz. Science, politics and economics of the Peruvian anchoveta fishery. Marine Policy, 3(3):201–210, July 1979. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X79900526.

Goncalves:1979:CMR

[Gon79] Maria Eduarda Gonçalves. Concepts of marine region and the new law of the sea. Marine Policy, 3(4):255–263, October 1979. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X79900162.

Gulland:1977:WFF

[Gul77a] J. A. Gulland. World fisheries and fish stocks. Marine Policy, 1 (3):179–189, July 1977. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0308597X77900252.

Gulland:1977:LRS

[Gul77b] John Gulland. Living resources of Southern Ocean: SCAR Symposium, Woods Hole, Mass, USA, 17–24 August 1976. Marine Policy, 1(1):79–80, January 1977. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X7790015X.

Gulland:1978:AOT

[Gul78] J. A. Gulland. Alternatives for the organization of tuna management. Marine Policy, 2(2):161–163, April 1978. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X78900118.

Guzhenko:1977:SMM

[Guz77] Timofei Guzhenko. Soviet merchant marine and world shipping. Marine Policy, 1(2):102–109, April 1977. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X77900458.

Hargreaves:1977:ISU

[Har77] P. M. Hargreaves. Information service in UK. Marine Policy, 1(4): 340–341, October 1977. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0308597X77900902.

Haslam:1977:HSB

[Has77] D. W. Haslam. Hydrographic services: Basic charting vital to all sea users. Marine Policy, 1(1):3–15, January 1977. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X77900033.

Heaver:1977:LEC

[Hea77] Trevor D. Heaver. Letters to the editor: Comecon shipping policy — international impacts and implications. Marine Policy, 1(4): 359–361, October 1977. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0308597X77901038.

Herke:1978:CAP

[Her78] William H. Herke. Is coastal aquaculture in the public interest? Marine Policy, 2(1):78–79, January 1978. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X78900696.

Hjul:1977:WFH

[Hju77] Peter Hjul. World fish harvest — problems and implications: Eleventh Session of the Committee on Fisheries (COFI) of the Food and Agriculture Organisation of the United Nations, Rome, Italy, 19–26 April 1977. Marine Policy, 1(3):263–264, July 1977. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X77900355.

Holt:1978:MFS

[Hol78] Sidney J. Holt. Managing fish stocks: the role of international organizations in the Mediterranean Area. Marine Policy, 2(2): 133–141, April 1978. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0308597X78900064.

Horrocks:1977:MPC

[Hor77] J. C. S. Horrocks. The 1973 marine pollution convention: Problems and solutions. Marine Policy, 1(1):52–60, January 1977. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X77900082.

Ion:1977:BRO

[Ion77] D. C. Ion. Book review: Optimal development of the North Sea’s oil fields: a study in divergent government and company interests and their reconciliation, by Peter R. Odell and Kenneth E. Rosing, Kogan Page, London, 1976, 189 pages, £30.00. Marine Policy, 1(3):266–267, July 1977. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X77900379.

Jacovides:1979:TAL

[Jac79] Andreas J. Jacovides. Three aspects of the law of the sea: Islands, delimitation and dispute settlement. Marine Policy, 3(4): 278–288, October 1979. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0308597X79900186.

Joseph:1978:LET

[JG78] James Joseph and J. A. Gulland. Letters to the editor: Tuna management — the debate continues. Marine Policy, 2(3):

239–240, July 1978. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0308597X78900441.

Jones:1978:MAU

[Jon78] Cecil Jones. Marine archaeology in the UK: a pattern of development. Marine Policy, 2(4):321–330, October 1978. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X78900210.

Joseph:1977:MHM

[Jos77] James Joseph. The management of highly migratory species: Some important concepts. Marine Policy, 1(4):275–288, October 1977. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X77900847.

Josephs:1978:JIU

[Jos78] Hilary K. Josephs. Japanese investment in the US fishing industry and its relation to the Two-Hundred-Mile Law. Marine Policy, 2(4):255–267, October 1978. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X78900179.

Josephs:1979:JIU

[Jos79] Hilary K. Josephs. Japanese investment in US fishing industry — update. Marine Policy, 3(2):164, April 1979. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X79900125.

Kaufmann:1977:UMS

[Kau77] Peter G. Kaufmann. Undersea medical society meeting: Toronto, Ontario, Canada, 13–16 May 1977. Marine Policy, 1(4):350– 352, October 1977. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0308597X77900975.

Kaufmann:1979:UMS

[Kau79] Peter G. Kaufmann. Undersea medical society meeting: Key Biscayne, Florida, USA, 29 May–2 June 1979. Marine Policy, 3(4): 324–325, October 1979. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0308597X79900277.

Keil:1977:UOO

[Kei77] Alfred A. H. Keil. The USA and the oceans: Opportunities for independence. Marine Policy, 1(1):16–25, January 1977. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X77900045.

Kelly:1978:FCM

[Kel78] John E. Kelly. The fishery conservation and management act of 1976: Organizational structure and conceptual framework. Marine Policy, 2(1):30–36, January 1978. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X78900581.

Kieth:1977:FCN

[Kie77] Kent M. Kieth. Floating cities: a new challenge for transnational law. Marine Policy, 1(3):190–204, July 1977. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X77900264.

Kimball:1977:FSU

[Kim77a] Lee Kimball. Fifth Session of the United Nations Conference on the Law of the Sea: New York, USA, 2 August–17 September 1976. Marine Policy, 1(1):80–82, January 1977. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X77900161.

Kimball:1977:SST

[Kim77b] Lee Kimball. Sixth session of the third United Nations conference on the law of the sea: New York, USA, 27 May–15 July 1977. Marine Policy, 1(4):344–346, October 1977. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X7790094X.

Kimball:1978:UGM

[Kim78] Lee Kimball. UNCLOS gains momentum. Marine Policy, 2(4): 334–337, October 1978. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0308597X78900246.

Kimball:1979:ESU

[Kim79a] Lee Kimball. Eighth session of the United Nations conference on the law of the sea: Geneva, Switzerland, 19 March–27 April 1979. Marine Policy, 3(4):317–320, October 1979. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X79900241.

Kimball:1979:UW

[Kim79b] Lee Kimball. UNCLOS — down to the wire? Marine Policy, 3 (2):156–158, April 1979. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0308597X79900071.

King:1979:IMH

[Kin79] Dennis M. King. International management of highly migratory species: Centralized versus decentralized economic decision making. Marine Policy, 3(4):264–277, October 1979. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X79900174.

Kingham:1979:CIO

[KM79] J. D. Kingham and D. M. McRae. Competent international organizations and the law of the sea. Marine Policy, 3(2):106– 132, April 1979. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0308597X79900046.

Knight:1978:MPU

[Kni78] H. Gary Knight. Management procedures in the US fishery conservation zone. Marine Policy, 2(1):22–29, January 1978. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X7890057X.

Kwasniewski:1977:SPC

[Kwa77] Klaus Kwasniewski. Shipping policy of the COMECON countries. Marine Policy, 1(2):110–117, April 1977. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X7790046X.

Larson:1979:SDL

[Lar79] David L. Larson. Security, disarmament and the law of the sea. Marine Policy, 3(1):40–58, January 1979. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X79900393.

Lawson:1977:NDD

[Law77a] Rowena M. Lawson. New directions in developing small-scale fisheries. Marine Policy, 1(1):45–51, January 1977. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X77900070.

Lawson:1977:WSS

[Law77b] Rowena M. Lawson. Workshop on small-scale fishery enterprise management, sponsored by the food and agriculture organisation and the Australian Development Assistance Agency: Colombo, Sri Lanka, 8 November–3 December 1976. Marine Policy, 1(2): 169–170, April 1977. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0308597X77900549.

Lennard:1977:BRJ

[Len77a] D. E. Lennard. Book review: Jane’s ocean technology 1976–77: edited by Robert L. Trillo, 679 pages, £25.00, 1976, Macdonald and Jane’s, London, UK. Marine Policy, 1(1):87–88, January 1977. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X77900227.

Lennard:1977:BRE

[Len77b] D. E. Lennard. Book review: The European offshore oil and gas yearbook 1976–77: edited by Ethel de Keyser, Kogan Page, London, 1976, 475 pages, £20.00. Marine Policy, 1(3):268– 269, July 1977. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0308597X77900392.

Lennard:1978:BRJ

[Len78] D. E. Lennard. Book review: Jane’s ocean technology 1978: edited by Robert L. Trillo, Third edition, MacDonald and Jane’s, London, 1978, 820 pages, £27.50. Marine Policy, 2(4):340– 341, October 1978. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0308597X78900271.

Lennard:1979:BRO

[Len79] D. E. Lennard. Book review: Offshore oil and gas yearbook 1978/79: UK and continental Europe: Kogan Page Ltd, London, 1978, 352 pp, £25. Marine Policy, 3(4):327–328, October 1979. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X79900307.

Loftas:1977:DO

[Lof77a] Tony Loftas. Decade of the oceans. Marine Policy, 1(1): 2, January 1977. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0308597X77900021.

Loftas:1977:NSO

[Lof77b] Tony Loftas. North sea oil: Who leads whom? Marine Policy, 1(3):178, July 1977. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0308597X77900240.

Loftas:1977:PPH

[Lof77c] Tony Loftas. Pity the poor herring. Marine Policy, 1(4): 274, October 1977. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0308597X77900835.

Loftas:1977:UPU

[Lof77d] Tony Loftas. UNCLOS: a parcel unpacked. Marine Policy, 1 (2):90, April 1977. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0308597X77900434.

Loftas:1978:K

[Lof78a] Tony Loftas. In for the krill. Marine Policy, 2(3):170, July 1978. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X78900362.

Loftas:1978:NCF

[Lof78b] Tony Loftas. New compensation fund: Closing the stable door …. Marine Policy, 2(4):254, October 1978. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X78900167.

Loftas:1978:RD

[Lof78c] Tony Loftas. Resources of discontent. Marine Policy, 2(1): 2, January 1978. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0308597X78900556.

Loftas:1978:TTI

[Lof78d] Tony Loftas. Third triennial international exhibition on the exploitation of the oceans, offshore techniques, shipbuilding, sea harbours and equipment: Bordeaux, France, 4–8 October 1977. Marine Policy, 2(1):68–71, January 1978. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X78900635.

Loftas:1978:UOM

[Lof78e] Tony Loftas. UNCLOS and ocean mining: Molehill or mountain? Marine Policy, 2(2):86, April 1978. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X78900015.

Loftas:1979:ELT

[Lof79a] Tony Loftas. Extended limits: a threat to small-scale fisheries. Marine Policy, 3(3):170, July 1979. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X79900496.

Loftas:1979:LSN

[Lof79b] Tony Loftas. Law of the sea: neglected issues. Marine Policy, 3 (3):243–245, July 1979. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0308597X79900587.

Loftas:1979:LSW

[Lof79c] Tony Loftas. Law of the sea: Who should do what? Marine Policy, 3(2):78, April 1979. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0308597X79900010.

Loftas:1979:RSR

[Lof79d] Tony Loftas. Resources — something in reserve? Marine Policy, 3(1):2, January 1979. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0308597X79900368.

Loftas:1979:SFD

[Lof79e] Tony Loftas. Safety first for divers. Marine Policy, 3(4): 254, October 1979. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0308597X79900150.

Lowe:1977:BRI

[Low77a] A. V. Lowe. Book review: The impact of flags of convenience: R. S. Doganis and B. N. Metaxas 137 pages, £6.00, 1976, Transport Studies Group Research Report 3, Polytechnic of Central London, London, UK. Marine Policy, 1(1):85–86, January 1977. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X77900197.

Lowe:1977:ILF

[Low77b] A. V. Lowe. International law and federal offshore lands disputes. Marine Policy, 1(4):311–317, October 1977. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X77900872.

Lowe:1978:TBC

[Low78] A. V. Lowe. Third biennial conference of the Greenwich Forum: Royal Naval College, Greenwich, London, UK, 4–7 October 1977. Marine Policy, 2(1):71–72, January 1978. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X78900647.

Macdonald:1979:IFI

[Mac79] R. D. S. Macdonald. Inshore fishing interests on the Atlantic coast: Their response to extended jurisdiction by Canada. Marine Policy, 3(3):171–189, July 1979. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X79900502.

Mankabady:1978:IRR

[Man78] Samir Mankabady. Insurance and reinsurance requirements for LNG. Marine Policy, 2(4):331–333, October 1978. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X78900222.

Mankabady:1979:FLC

[Man79] Samir Mankabady. Financing LNG carriers. Marine Policy, 3 (3):225–231, July 1979. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0308597X7990054X.

Marjoram:1979:NSM

[Mar79] Tony Marjoram. Nodules and Scandinavia — mining opportunities. Marine Policy, 3(4):313–315, October 1979. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X79900228.

Mason:1977:IPN

[Mas77] C. M. Mason. International politics of the North Sea. Marine Policy, 1(1):73–74, January 1977. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X77900124.

Mathers:1978:DLU

[Mat78] Nigel F. Mathers. Diving legislation in the UK. Marine Policy, 2(3):227–233, July 1978. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X78900404.

McKnight:1979:FRO

[McK79] Annena Billings McKnight. Federal role in oceans development. Marine Policy, 3(2):158–159, April 1979. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X79900083.

Meirowitz:1978:BRE

[Mei78] Mark Meirowitz. Book review: Economic impacts of extended fisheries jurisdiction: edited by Lee G. Anderson, Ann Arbor Science, Ann Arbor, Michigan, 1977, 428 pages, £16.90, $22.80. Marine Policy, 2(2):165–167, April 1978. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X78900131.

Metaxas:1978:BRF

[Met78] B. N. Metaxas. Book review: Federal port policy in the United States: by Henry S. Marcus, James E. Short, John C. Kuypers, and Paul O. Roberts, MIT Press, Cambridge, Mass, 1977, 371 pages, £15.00. Marine Policy, 2(3):249–250, July 1978. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X78900520.

Metaxas:1979:BRM

[Met79] B. N. Metaxas. Book review: Middle East liner shipping: an economic analysis of traffic, services, ports and future prospects. H. P. Drewry (Shipping Consultants) Ltd Publications, 34 Brook Street, London W1Y2LL, 1978, 144 pp. Marine Policy, 3(3): 246–247, July 1979. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0308597X79900605.

Mitchell:1977:RAP

[Mit77] Barbara Mitchell. Resources in Antarctica: Potential for conflict. Marine Policy, 1(2):91–101, April 1977. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X77900446.

Morris:1978:BRM

[Mor78] R. J. Morris. Book review: Marine pollution: edited by R. Johnston, Academic Press. London and New York, 1977, 729 pages, £23.00, & 51.00. Marine Policy, 2(1):75–76, January 1978. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X78900660.

Mackay:1978:OFC

[MP78] G. A. Mackay and D. W. Pearce. Oil and fishing — conflict in the North Sea. Marine Policy, 2(4):333–334, October 1978. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X78900234.

Nash:1979:BRA

[Nas79] Colin E. Nash. Book review: Aquaculture in the United States: Constraints and opportunities: the National Research Council, under the auspices of the National Academy of Sciences, Washington, DC 20418, 1978. 123pp. Marine Policy, 3(3):248–250, July 1979. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X79900629.

OConnell:1977:BRN

[O’C77] D. P. O’Connell. Book review: Navies and foreign policy: K. Booth, Croom Helm, London, 1977, 294 pages, £9.95. Marine Policy, 1(4):352–353, October 1977. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X77900987.

OConnell:1978:BRN

[O’C78] D. P. O’Connell. Book review: New directions in the law of the sea: Vols 5 and 6, compiled and edited by R. Churchill, M. Nordquist and S. Houston Lay, Oceans and the British Institute of International and Comparative Law, London, 1977, £16.50 per volume. Marine Policy, 2(4):339–340, October 1978. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X7890026X.

OConnell:1979:BRT

[O’C79] D. P. O’Connell. Book review: Technical change and British naval policy 1860–1939: edited by Bryan Ranft, Hodder and Stoughton, London, 1977, 178pp. Marine Policy, 3(1):73–74, January 1979. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X79900460.

Ogley:1977:BRN

[Ogl77] Roderick C. Ogley. Book review: The new international economic order and the law of the sea: Arvid Pardo and Elisabeth Mann Borgese, 223 pages, $10.00, 1976, Occasional Paper No 4, International Ocean Institute, Malta. Marine Policy, 1(1): 83–85, January 1977. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0308597X77900185.

Ogley:1979:BRE

[Ogl79] Roderick C. Ogley. Book review: The enterprises: by Elisabeth Mann Borgese, International Ocean Institute Occasional Papers No 6, University of Malta Press, Malta, November 1978, 23 pp, no price given. Marine Policy, 3(3):247–248, July 1979. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X79900617.

Palacio:1979:APL

[Pal79] Francisco J. Palacio. Aquaculture policies in Latin America. Marine Policy, 3(2):99–105, April 1979. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X79900034.

Pantelidis:1979:GSA

[Pan79] Evangelos T. Pantelidis. Greek shipping and the accession of Greece to the EEC. Marine Policy, 3(1):3–19, January 1979. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X7990037X.

Parrish:1979:FRI

[Par79] B. B. Parrish. The future role of ICES in the light of changes in fisheries jurisdiction. Marine Policy, 3(3):232–238, July 1979. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X79900551.

Portmann:1977:BRS

[Por77] J. E. Portmann. Book review: Strategies for marine pollution monitoring: Edward D. Goldberg, John Wiley and Sons, New York, London, Sydney, Toronto, 1976, 310 pages. £17.05. $27.50. Marine Policy, 1(3):265–266, July 1977. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X77900367.

Preston:1978:RWS

[Pre78] A. Preston. Radioactive wastes in the seabed?: Oceanus, vol 20, no 1, 1977. Marine Policy, 2(1):76–77, January 1978. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X78900672.

Pritchard:1977:BRC

[Pri77a] S. Z. Pritchard. Book review: The control of oil pollution on the sea and inland waters: edited by J. Wardley-Smith, Graham and Trotman, London, 1976, 251 pages, £9.75. Marine Policy, 1 (3):267–268, July 1977. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0308597X77900380.

Pritchard:1977:BRI

[Pri77b] T. Pritchard. Book review: An International Conference on Marine Parks and reserves (Tokyo, Japan, May 1975): 132 pages, $10.00, 1976, International Union for Conservation of Nature and Natural Resources, Merges, Switzerland. Marine Policy, 1(1): 86–87, January 1977. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0308597X77900215.

Pritchard:1978:NSS

[Pri78] S. Z. Pritchard. The North Sea as a special area. Marine Policy, 2(1):65–67, January 1978. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X78900611.

Pullin:1977:AES

[Pul77a] R. S. V. Pullin. Aquaculture in Europe: the Second European Fish Farming Congress, sponsored by the European Mariculture Society, Fish Farming International, and the National Farmers’ Union of the UK, London, UK, 30 November–1 December 1976. Marine Policy, 1(3):259–261, July 1977. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X77900331.

Pullin:1977:BRF

[Pul77b] R. S. V. Pullin. Book review: Farming the edge of the sea: E. S. Iversen, Fishing News Books, Farnham, Surrey, 1976, second edition, 436 pages, £13.50. Marine Policy, 1(4):356, October 1977. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X77901014.

Pullin:1977:CAE

[Pul77c] R. S. V. Pullin. Coastal aquaculture in Europe. Marine Policy, 1(3):205–214, July 1977. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X77900276.

Pullin:1979:BRS

[Pul79a] R. S. V. Pullin. Book review: Salmon and trout farming in Norway: by David J. Edwards, Fishing News Books Limited, Farnham, Surrey, UK, 1978, 195 pp. £9.00. Marine Policy, 3(2): 163–164, April 1979. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0308597X79900113.

Pullin:1979:CCE

[Pul79b] R. S. V. Pullin. Conflicts and cooperation in European fisheries. Marine Policy, 3(2):160–161, April 1979. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X79900095.

Ramster:1977:DCR

[Ram77] J. W. Ramster. Development of cooperative research in the North Sea: the origins, planning and philosophy of JONSDAP 76. Marine Policy, 1(4):318–325, October 1977. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X77900884.

Rawson:1978:OP

[Raw78] Kenneth Rawson. Offshore policeman. Marine Policy, 2(1):37– 45, January 1978. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0308597X78900593.

Reynolds:1977:ELS

[Rey77] Paul D. Reynolds. The EEC and the law of the sea. Marine Policy, 1(2):118–131, April 1977. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X77900471.

Richards:1979:FPC

[Ric79a] Robert A. C. Richards. Fisheries policy for the Canary Isles. Marine Policy, 3(3):239–240, July 1979. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X79900563.

Richards:1979:SFF

[Ric79b] Robert A. C. Richards. Spain’s foreign fishing policy. Marine Policy, 3(4):312–313, October 1979. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X79900216.

Richards:1979:SAS

[Ric79c] Robert A. C. Richards. Spanish adaptation to shipyards crisis. Marine Policy, 3(1):67–69, January 1979. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X79900423.

Rizzi:1978:ETM

[Riz78] Daniele Rizzi. The EEC treaty and the merchant fleets of the member states. Marine Policy, 2(4):268–274, October 1978. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0308597X78900180.

Roberts:1979:IMR

[RS79] Peter W. Roberts and Timothy Shaw. Impacts and management of resource exploitation. Marine Policy, 3(1):66–67, January 1979. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X79900411.

Russell:1978:BRE

[Rus78] Clifford S. Russell. Book review: An environmental appraisal of the Severn Barrage: edited by T. L. Shaw, Department of Civil Engineering, University of Bristol, 1977, 2nd ed, 155 pages, £3.00. Marine Policy, 2(1):72–74, January 1978. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X78900659.

Russell:1977:HSE

[RV77] Clifford S. Russell and William J. Vaughan. Hydrographic services — essential improvements. Marine Policy, 1(2):174–176, April 1977. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X77900586.

Sakiyama:1979:PPA

[Sak79] Teruji Sakiyama. Policies in pollution, aquaculture and coastal management in Japan. Marine Policy, 3(1):20–39, January 1979. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X79900381.

Saliba:1978:PMC

[Sal78] Louis J. Saliba. Protecting the Mediterranean: Coordinating regional action. Marine Policy, 2(3):171–180, July 1978. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X78900374.

Scott:1978:TSA

[Sco78] Desmond P. D. Scott. Tenth session of the assembly of the Intergovernmental Oceanographic Commission: Paris, France, 27 October–10 November 1977. Marine Policy, 2(2):158–159, April 1978. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X7890009X.

Sebek:1979:UPC

[Seb79] Viktor Sebek. UK pollution control — Whitehall v. Parliament. Marine Policy, 3(3):240–242, July 1979. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X79900575.

Shaw:1977:PTE

[Sha77a] T. L. Shaw. A policy for tidal energy. Marine Policy, 1(1): 61–69, January 1977. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0308597X77900094.

Shaw:1977:TEI

[Sha77b] T. L. Shaw. Tidal energy — ideal source. Marine Policy, 1(3): 270–271, July 1977. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0308597X77900410.

Shackleton:1979:ESC

[Sha79a] Michael Shackleton. Europe and the sea: the case for and against a new international regime for the North Sea and its approaches, the Greenwich forum international conference: Royal Naval College, Greenwich, UK, 2–4 May 1979. Marine Policy, 3(4):320– 322, October 1979. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0308597X79900253.

Shaw:1979:BRM

[Sha79b] T. L. Shaw. Book review: Marine outfall systems: Planning, design and construction: Robert A. Grace, Prentice-Hall International, Englewood Cliffs, New Jersey, USA, 600 pp, £25.55. Marine Policy, 3(4):326, October 1979. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X79900289.

Shepherd:1978:BRA

[She78] C. J. Shepherd. Book review: Aquaculture for the developing countries: by Frederick W. Bell and E. Ray Canterbery, Ballinger, Cambridge, Mass, 1977, 267 pp, £13.40, $22.00. Marine Policy, 2(3):245–247, July 1978. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X78900490.

Siemens:1977:FP

[Sie77] Abe H. Siemens. Fighting pollution. Marine Policy, 1(1):70– 71, January 1977. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0308597X77900100.

Simpson:1978:RRF

[Sim78] A. C. Simpson. The role of research in fisheries development. Marine Policy, 2(3):212–226, July 1978. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X78900398.

Snow:1978:UCP

[Sno78] Beverly C. Snow. US Coastal Plains Marine Center. Marine Policy, 2(3):234–235, July 1978. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X78900416.

Stansell:1977:ODW

[Sta77] John Stansell. Oil in deeper waters: Second International Offshore Exhibition and Conference, organised by The Financial Times, Birmingham, UK. 7–10 December 1976. Marine Policy, 1(3): 261–263, July 1977. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0308597X77900343.

Stanford:1978:OPO

[Sta78] Gerald B. Stanford. Offshore policemen: operational and legal considerations. Marine Policy, 2(4):343–344, October 1978. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0308597X78900325.

Stokes:1979:LFE

[Sto79] Robert L. Stokes. Limitation of fishing effort: an economic analysis of options. Marine Policy, 3(4):289–301, October 1979. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0308597X79900198.

Sturmey:1979:DCC

[Stu79] S. G. Sturmey. The development of the code of conduct for liner conferences. Marine Policy, 3(2):133–148, April 1979. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X79900058.

Tetley:1978:BRC

[Tet78] Michele Mason Tetley. Book review: Coastal ecosystem management: a technical manual for the conservation of coastal zone resources: by John Clark, John Wiley and Sons, London and New York, 7977, 925 pp, £28.50, $48.50. Marine Policy, 2 (3):242, July 1978. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0308597X78900465.

Tambs-Lyche:1978:MFS

[TL78] Hans Tambs-Lyche. Monitoring fish stocks: the role of ICES in the North-East Atlantic. Marine Policy, 2(2):127–132, April 1978. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X78900052.

Tomczak:1979:CE

[Tom79] Matthias Tomczak. The ClNECA experience. Marine Policy, 3 (1):59–65, January 1979. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0308597X7990040X.

Valencia:1978:SCS

[Val78] Mark J. Valencia. The South China Sea: Prospects for marine regionalism. Marine Policy, 2(2):87–104, April 1978. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X78900027.

Watt:1977:BRS

[Wat77a] D. C. Watt. Book review: Sea power and the law of the sea: Mark W. Janis Lexington Books, distributed in UK by Teakfield Ltd, Farnborough, Hants, 1977, 99 pages, £6.50. Marine Policy, 1(4):354–355, October 1977. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X77901002.

Watt:1977:CSF

[Wat77b] D. C. Watt. Conference on sea floor development — Moving into deeper waters, organized by the Royal Society: London, UK, 1–2 June 1977. Marine Policy, 1(4):347–348, October 1977. CODEN ???? ISSN 0308-597X (print), 1872-9460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0308597X77900951.

Watt:1978:BRN

[Wat78] D. C. Watt. Book review: New strategic factors in the North Atlantic: edited by Christoph Bertram and Johan J. Holst, IPC Science and Technology Press, Guildford, and Universitetsforlaget, Oslo, 1977, 194 pp, £11.50, $25.00. Marine Policy, 2(3): 244–245, July 1978. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0308597X78900489.

Watt:1979:BRM

[Wat79a] D. C. Watt. Book review: Marine policy: a comparative approach: by John King Gamble Jr Lexington Books, Lexington, MA, and Famborough. Hants, 1977, 147 pp. Marine Policy, 3(1): 72–73, January 1979. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0308597X79900459.

Watt:1979:BRR

[Wat79b] D. C. Watt. Book review: Resource management at the international level: the case of the North Pacific: by Oran J. Young, Francis Pinter, London and New York, 1977, 252 pp. Marine Policy, 3(1):72–73, January 1979. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X79900447.

Watt:1979:BRV

[Wat79c] D. C. Watt. Book review: Volkerrechtliche und politische Auswirkungen der dritten Seerechtskon Ferenz Vereinten Nationen auf Nord und Ostsee: by Renate Platzöder, Publications of the Stiftung Wissenschaft und Politik Forschungsinstitut für internationalen Politik und Sicherheit, Nr SWP-S 266, Hans Eggenberg, Ebenhausen bei München, October 1978, 211 pp. Marine Policy, 3(4):328–329, October 1979. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X79900319.

Watt:1979:FSE

[Wat79d] Donald Cameron Watt. First steps in the enclosure of the oceans: the origins of Truman’s proclamation on the resources of the continental shelf, 28 September 1945. Marine Policy, 3 (3):211–224, July 1979. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0308597X79900538.

Wyner:1977:PMC

[WMCS77] Alan J. Wyner, John E. Moore, and Biliana Cicin-Sain. Politics and management of the California abalone fishery. Marine Policy, 1(4):326–339, October 1977. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X77900896.

Woodliffe:1978:FCF

[Woo78] J. C. Woodliffe. Floating cities: further thoughts on their legal status. Marine Policy, 2(1):79–81, January 1978. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X78900702.

Yonge:1977:JOA

[Yon77] Sir Maurice Yonge. Joint oceanographic assembly 1976: Edinburgh, UK, 13–24 September 1976. Marine Policy, 1(2):168– 169, April 1977. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0308597X77900537.

Young:1977:BRM

[You77] Elizabeth Young. Book review: Maritime policy and the coastal community — The impact of the law of the sea: edited by Douglas M. Johnston, 338 pages, £9.95, 1976, Croom Helm, London, UK. Marine Policy, 1(1):82–83, January 1977. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X77900173.

Young:1978:BRPb

[You78a] Elizabeth Young. Book review: Pacific basin enterprise and the changing law of the sea: Jeremiah J. Sullivan, Pacific Rim Research Series No 2, Lexington Books, Lexington, MA, 1977, 219 pages. Marine Policy, 2(4):341–342, October 1978. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X78900295.

Young:1978:BRPa

[You78b] Elizabeth Young. Book review: The Pacific quest: Endel-Jakob Kolde, Pacific Rim Research Series No 1, Lexington Books, Lexington, MA, 1976, 162 pages. Marine Policy, 2(4):341– 342, October 1978. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0308597X78900283.

Young:1979:MAL

[You79] Elizabeth Young. Military activities and law of the sea. Marine Policy, 3(3):251, July 1979. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X79900642.